OpenStudy (anonymous):

A piece of wire 10m long is cut into two pieces. One piece is bent into a square and the other is bent into an equilateral triangle. How should the wire be cut so that the total are enclosed is (a) a maximum? (B) Is a minimum? Help please

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

The total volume it going to be"\[Volue_{total} = Volume_{\square} +Volume_{\triangle}\]

OpenStudy (anonymous):

the constraints on these volume is the perimeter. Set the sides of one of the object to a variable, say, x. then the perimeter will be related through the length of original wire: Let x = length of side on square Perimeter of square: = 4x Perimeter of Tri = 10m - 4x Side of Triangle = (10-4x)/3

OpenStudy (anonymous):

Now express the Volume (the quantity you wish to maximize) as a function of the variable: \[Volume_{total} = volume_{\square}+Volume_{\triangle}\] \[Volume_{total} =(length*width)_{\square}+\frac{1}{2}(base*height)_{\triangle}\] \[Volume_{total} = (x*x)+\frac{1}{2}(\frac{10-4x}{2})(\sqrt3(\frac{10-4x}{2}))\]

OpenStudy (anonymous):

Why are you setting it equal to the volume?

OpenStudy (anonymous):

Because the question asks for you to solve for the maximum/minimum area enclosed depending on how you cut the wire. To solve Max/min question you need an equation that you can take the derivative of, in this case Volume with respect to length

OpenStudy (anonymous):

derivatives allow you to solve max/min because when the derivative is zero, that represents a hump (max/min) in the original function

OpenStudy (anonymous):

thanks a lot, I get it now

OpenStudy (anonymous):

hahah..i see. I've written volume when I've meant Area throughout the whole question. please substitute Area

OpenStudy (anonymous):

One more question, how did you get the height of the triangle?

OpenStudy (anonymous):

Ya, I figured!!

OpenStudy (anonymous):

special triangle (30 degrees/60 degrees) or straight pythagorean relation